Show that low-frequency limit of Planck’s Law reduces to the Rayleigh-Jeans Law


and in the high-frequency limit reduces to Wien’s Law

1
Expert's answer
2022-03-13T18:47:27-0400

The Planck’s Law says

"B(\nu,T)=\frac{2h\nu^3}{c^2}\frac{1}{e^{h\nu/kT}-1}"

At low-frequency

"e^{h\nu/kT}-1=1+h\nu/kT-1=h\nu/kT"

Hence

"B(\nu,T)=\frac{2h\nu^3}{c^2}\frac{1}{h\nu/kT}=\frac{2\nu^2}{c^2}kT"

At high-frequency

"e^{h\nu/kT}-1=e^{h\nu/kT}"

Hence

"B(\nu,T)=\frac{2h\nu^3}{c^2}\frac{1}{e^{h\nu/kT}}=\frac{2h\nu^3}{c^2}e^{-h\nu/kT}"